OpenStudy (anonymous):

derivate 2sin(X)-cos(x)-cot(x)

OpenStudy (anonymous):

d/dx 2sin(x) -cos(x) -cot(x)= 2cos(x) -(-sin(x)) -(-1/(sin(x))^2) 2cos(x) + sin(x) + 1/(((sin(x))^2)

OpenStudy (anonymous):

i thought derivate of cot is -csc^2

OpenStudy (anonymous):

correct and csc(x) = 1/sin(x), so csc^2 = 1/sin(x)^2.

OpenStudy (anonymous):

thank you i see it now :=)

OpenStudy (anonymous):

good ; )